OpenStudy (anonymous):

If the unit price for cantaloupe is $0.65 per pound and Sally pays $1.95 for a cantaloupe. how many pounds does her cantaloupe weigh. 2 pounds 3 pounds 4 pounds 5 pounds

OpenStudy (anonymous):

\[1.95\div .65\]

OpenStudy (anonymous):

3 pounds? @satellite73

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thanks will you help with a couple more I'll fan. you pls pls pls pls pls pls pls!!!!! @satellite73

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

Omg thanks ok.

OpenStudy (anonymous):

Which pair of ratios can form a true proportion? 1/2 6/10 6/8 5/20 3/5 6/10 3/9 7/80

OpenStudy (anonymous):

you could try cross multiplying, or reduce each fraction, or just go with C because \[\frac{3}{5}\times \frac{2}{2}=\frac{6}{10}\]

OpenStudy (anonymous):

Solve the proportion Y/22=3/11. y=6 y=11 y=15 y=66

OpenStudy (anonymous):

\[\frac{y}{22}=\frac{3}{11}\\ y=\frac{3\times 22}{11}\] cancel first multiply last

OpenStudy (anonymous):

let me know when you get \(y=6\)

OpenStudy (anonymous):

The length of a kitten on a scale drawing is 8 cm. If the scale is 1 cm:6 cm, what is the actual length of the kitten?

OpenStudy (anonymous):

yes I got 6 why?

OpenStudy (anonymous):

because that is right

OpenStudy (anonymous):

Oh thanks.

OpenStudy (anonymous):

The length of a kitten on a scale drawing is 8 cm. If the scale is 1 cm:6 cm, what is the actual length of the kitten? \[8\times 6\]

OpenStudy (anonymous):

The distance between spice island and Kay island is 3.4 inches on a map. If the scale on the map is 1 inch =25 miles, what is the actual distance between the two islands. 28.4 75 85 850

OpenStudy (anonymous):

\[3.4\times 25\]

OpenStudy (anonymous):

Solve the proportion 4/8=w/32.

OpenStudy (anonymous):

options for that one above is 2 7 8 16

OpenStudy (anonymous):

\[\frac{4}{8}=\frac{1}{2}\] so \[\frac{1}{2}=\frac{w}{32}\iff w=\frac{32}{2}\]

OpenStudy (anonymous):

I have 10 more questions is that too much if so thanks for the help so far

OpenStudy (anonymous):

go ahead we can see what we can do

OpenStudy (anonymous):

Thanks. Ok Which of the following are equivalent to 45%select that all apply 0.045 0.45 9/10 9/20

OpenStudy (anonymous):

\[0.45=45\%\]

OpenStudy (anonymous):

that is all

OpenStudy (anonymous):

oh no sorry \(\frac{9}{20}=\frac{45}{100}=45\%\) so there are two of them

OpenStudy (anonymous):

Select the answer that correctly order the set of the numbers from greatest to least. 1/8, 36%, 0.85, 0.078

OpenStudy (anonymous):

oh you want "greatest to least" right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[.85,36\%,\frac{1}{8},0.078\]

OpenStudy (anonymous):

if you convert them all to decimals it is easy to see it

OpenStudy (anonymous):

If 14 out of 20 people like chocolate milk, what percent of the people do not like chocolate milk? 5% 30% 40% 70%

OpenStudy (anonymous):

\[\frac{6}{20}=\frac{30}{100}=30\%\]

OpenStudy (anonymous):

For greatest to least question what is it.

OpenStudy (anonymous):

answer above

OpenStudy (anonymous):

Oh. sorry.

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Sally is photocopying a picture. The original picture is 4 inches wide and 6 inches long. The new copy is 16 inches wide. How long is the new copy? 12 in. 8 in. 24 in. 36 in.

OpenStudy (anonymous):

\[6\times 4\]

OpenStudy (anonymous):

What is 26% of 125

OpenStudy (anonymous):

\[0.26\times 125\] and a calculator

OpenStudy (anonymous):

What is the best estimate of 26% of 81 6 20 45 58

OpenStudy (anonymous):

i would say it is like one quarter of 80, which is 20

OpenStudy (anonymous):

If you were to sketch a circle graph for data shown table shown below , about what fraction of the circle would represent comedy Types Vote Comedy 45 adventure 56 mystery 39

OpenStudy (anonymous):

Options for post above 1/2 1/3 1/4 1/5

OpenStudy (anonymous):

\[45+56+39=140\] so \[\frac{45}{140}\]

OpenStudy (anonymous):

i guess you are supposed to day \(\frac{1}{3}\)

OpenStudy (anonymous):

3 more I promise Lol. If the original bill is $15.50, what will be the final bill after paying a 15% tip? round to the nearest cent.

OpenStudy (anonymous):

\[15.5\times .15\]

OpenStudy (anonymous):

What is the amount of sales tax owed on a $350 laptop computer if the tax rate is 6%

OpenStudy (anonymous):

\[350\times .06\]

OpenStudy (anonymous):

the tip one i got wrong you want the total amount

OpenStudy (anonymous):

that would be about \(\$17.83\)

OpenStudy (anonymous):

Find the amount of the discount on a $120 bicycle that is on sale for 25% off. $25 $30 $73 $90

OpenStudy (anonymous):

\[120\times .25\]

OpenStudy (anonymous):

of one fourth of 120, which is 30

OpenStudy (anonymous):

And thats it thank sooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooo much ill fan i have 4 more but ill ask someone else u helped enough

OpenStudy (anonymous):

are they easy?

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

go ahead lets knock them out

OpenStudy (anonymous):

Omg really thanks sooo much.

OpenStudy (anonymous):

with these you have to explain.

OpenStudy (anonymous):

Find the unit price for each option shown below. Round to the nearest cent when necessary. Indicate which option is better to buy. Option 1: 10 candy bars for $6.75 option 2: 12 candy bars for $7.25

OpenStudy (anonymous):

price per candy bar is \[\frac{$}{\text{number of bars}}\]

OpenStudy (anonymous):

compare \[\frac{6.75}{10}\] with \[\frac{7.25}{12}\] and see which one is cheapest

OpenStudy (anonymous):

Option 1?

OpenStudy (anonymous):

\[6.75\div 10=.675\\ 7.25\div 12=.604\] so the second is cheapest

OpenStudy (anonymous):

Oh ok I see.

OpenStudy (anonymous):

Calculate the final price you would pay to buy jacket that originally sells for $65, but is discounted by 40%; you also have to pay 5% sales tax

OpenStudy (anonymous):

\[65\times .6\times 1.05\] and a calculator

OpenStudy (anonymous):

what does 'and a calculator' mean

OpenStudy (anonymous):

use a calculator to compute it

OpenStudy (anonymous):

does it mean like to calculate that to get your answer or...

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so the answer would be 58.5 right?

OpenStudy (anonymous):

not what i get no

OpenStudy (anonymous):

www.wolframalpha.com/input/?i=65*.6*1.05

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

i got 40.95 when i did it

OpenStudy (anonymous):

so 65 times .6 times 1.05 is 40.95??

OpenStudy (anonymous):

according to my calculator yes

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Katy was picking berries. She picked 270 in 18 minutes. How many will she have picked in 25 minutes.

OpenStudy (anonymous):

\[\frac{270}{18}=\frac{x}{25}\] \[x=\frac{25\times 270}{18}\]

OpenStudy (anonymous):

solve the proportion. 16/50=x over156.25

OpenStudy (anonymous):

\[x=\frac{156.25\times 16}{50}\]

OpenStudy (anonymous):

all of that equals 50

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

Can we break its time for dinner?

OpenStudy (anonymous):

lol go eat dear ttyl

OpenStudy (anonymous):

Turns out the the chicken isn't done yet lol. Use the circle graph shown below to answer the question 42% =football 25% =soccer 33% =baseball If 210 people said football was their favorite sport to watch how many people were surveyed?

OpenStudy (anonymous):

@iambatman can u help with the problem above above the sports